ACTIVITY 1

Consider the following equation:

Ca(OH)2(s) + 2 HCl(aq) —> CaCl2(aq) + 2 H2O(l)

a) How many liters of 0.100 M HCl is required to completely react with 5.00 grams of calcium hydroxide?

b) If 15.0 grams of calcium hydroxide is combined with 75.0 mL of 0.500 M HCl, how many grams of calcium chloride would be formed?


1
Expert's answer
2022-04-04T17:05:07-0400

Ca(OH)2(s) + 2 HCl(aq) —> CaCl2(aq) + 2 H2O(l)


1) According to the equation, n (HCl) = 2 x n (Ca(OH)2)

M (Ca(OH)2) = 74 g/mol

n (Ca(OH)2) = m / M = 5 / 74 = 0.07 mol

n (HCl) = 2 x n (Ca(OH)2) = 2 x 0.07 = 0.14 mol

CM (HCl) = n / V

V (HCl) = n / CM (HCl) = 0.14 / 0.1 = 1.4 L


2) n (Ca(OH)2) = m / M = 15 / 74 = 0.2 mol

n (HCl) = CM x V = 0.500 x 0.0750 = 0.038 mol

HCl is the limiting reactant.

n (CaCl2) = 1/2 x n (HCl) = 0.038 / 2 = 0.019 mol

n = m /M

M (CaCl2) = 111 g/mol

m (CaCl2) = n x M = 0.019 x 111 = 2.11 g